# Reflectance Prediction-based Knowledge Distillation for Robust 3D Object Detection in Compressed Point Clouds

### v1

@@ Abstract
Regarding intelligent transportation systems for vehicle networking, low-bitrate transmission via lossy point cloud compression is vital for facilitating real-time collaborative perception among vehicles with restricted bandwidth. In existing compression transmission systems, the sender lossily compresses point coordinates and reflectance to generate a transmission code stream, which faces transmission burdens from reflectance encoding and limited detection robustness due to information loss. To address these issues, this paper proposes a 3D object detection framework with reflectance prediction-based knowledge distillation (RPKD). We compress point coordinates while discarding reflectance during low-bitrate transmission, and feed the decoded non-reflectance compressed point clouds into a student detector. The discarded reflectance is then reconstructed by a geometry-based reflectance prediction (RP) module within the student detector for precise detection. A teacher detector with the same structure as student detector is designed for performing reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD) from raw to compressed point clouds. Our RPKD framework jointly trains detectors on both raw and compressed point clouds to improve the student detector’s robustness. Experimental results on the KITTI dataset and Waymo Open Dataset demonstrate that our method can boost detection accuracy for compressed point clouds across multiple code rates. Notably, at a low code rate of 2.146 Bpp on the KITTI dataset, our RPKD-PV achieves the highest mAP of 73.6, outperforming existing detection methods with the PV-RCNN baseline.
Index Terms:
Compressed Point Clouds, 3D Object Detection, Knowledge Distillation, Reflectance Prediction.I Introduction
@@
LiDAR point clouds, composed of sparse points with both geometric (coordinates) and attribute (reflectance) information about object surfaces, are vital for predicting the categories, positions, and bounding boxes of 3D objects. In cooperative perception among networked vehicles, local vehicles (receivers) need to rapidly acquire LiDAR point clouds captured by neighboring vehicles (senders) to enable an expanded perception range and accurate object detection [1, 2, 3, 4]. Raw point clouds are extremely voluminous and require a high transmission bandwidth, making it difficult to realize real-time cooperative perception. To this end, point cloud compression reduces the amount of transmitted data, improving data acquisition efficiency at the receiver. However, compressed point clouds generated by existing methods often lose some geometric coordinates and reflectance information, leading to suboptimal performance in 3D object detection. Therefore, enhancing the 3D detector’s robustness for compressed point clouds is of great research significance.
@@
Existing 3D object detection methods in compressed point clouds can be classified into two categories based on their reflectance processing approaches: 3D object detection with or without reflectance encoding. For 3D object detection with reflectance encoding, the geometry-based point cloud compression (G-PCC) [5] separately compresses the coordinates and reflectance of voxelized point clouds to form geometry and attribute bitstreams for reconstruction and detection tasks, as shown in Fig. 1(a). Building on this, Unicorn [6, 7] introduces a learning-based compression framework that utilizes multiscale sparse tensors from single-frame voxelized point clouds and lower-scale priors from multi-frame point clouds to improve compression and reconstruction performance. However, these methods face a significant burden from the attribute bitstream during networking transmission, limiting transmission efficiency. To solve this issue, in 3D object detection without reflectance encoding, existing compression methods based on voxelized point clouds and deep learning [8, 9, 10, 11, 12, 13] typically focus on compressing geometric coordinates while discarding reflectance to elevate transmission efficiency, as presented in Fig. 1(b). As an illustration, PCC-S [9] employs a geometric perception module to improve the learning of sibling and ancestor contexts in an octree structure for excellent geometric coordinates compression and reconstruction, during which the decoded data is termed non-reflectance compressed point clouds. However, lossy compression by these methods causes substantial information loss in low-bitrate transmission, severely impacting 3D object detection accuracy. To address these challenges, this paper aims to improve the transmission efficiency of lossy compressed point clouds and develop robust compressed point-cloud detectors, supported by a series of detailed studies.
Fig. 1(c) shows a baseline of our method. It takes accurate object predictions from raw point clouds as teacher knowledge for distillation during detection training. In this setup, the teacher model transfers its knowledge to compressed point-cloud detectors through response-based detection knowledge distillation (DKD). As shown in Fig. 2, SparseKD-PV outperforms PV-RCNN by an average of 1.11% in mAP across multiple data types, whereas it still exhibits a significant detection performance gap compared to PV-RCNN’s mAP of 75.56 on raw data. In general, LiDAR point cloud reflectance provides crucial cues for identifying 3D objects, thus improving the reliability of 3D object detection. In SparseKD-PV, the decision to discard reflectance information to enhance transmission efficiency is a key factor contributing to the suboptimal performance. While methods like Unicorn [7] reconstruct reflectance via a residual-based attribute bitstream using a DNN network, and PCC-S [9] does not transmit the attribute bitstream but needs alternative strategies to recover reflectance, we argue that relying solely on non-reflectance compressed point clouds for detection misses important reflectance features. To address this limitation, we propose a dedicated reflectance prediction (RP) module at the receiver, which utilizes the geometric information of compressed point clouds and reflectance priors from raw point clouds. Additionally, to further enhance the accuracy of reflectance prediction, we introduce response-based reflectance knowledge distillation (RKD), which effectively guides the RP module during learning process, improving detection performance while maintaining efficient data transmission.
@@
Incorporating the previous analysis, we propose a 3D object detection framework with reflectance prediction-based knowledge distillation (RPKD) to improve detection accuracy for non-reflectance compressed point clouds, as depicted in Fig. 1(d). Since compressed and raw point clouds lack direct point-to-point correspondence, we design a reflectance cross-match (RCM) module to assign reflectance labels from raw point-cloud voxels to compressed points. To construct suitable knowledge for reflectance distillation, we introduce a comparable reflectance inter-match (RIM) module to generate reflectance labels for the raw point-cloud detector. Besides, the proposed reflectance prediction (RP) module comprehensively extracts geometric features of voxelized point clouds, achieving reflectance reconstruction at the receiver. Through our reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD), rich teacher knowledge is transferred from raw to compressed point clouds for more precise reflectance reconstruction and robust 3D object detection. As shown in Fig. 2, RPKD-PV significantly increases the mAP values across multiple data types, which demonstrates the effectiveness of our RPKD framework for non-reflectance compressed point clouds.
@@
We introduce two response-based distillation constraints, namely RKD and DKD, to transfer supervisory knowledge from raw to compressed point-cloud detectors, further enhancing detection robustness for compressed point clouds.
The remainder of the paper is organized as follows. Section II reviews existing studies in terms of 3D object detection, point cloud compression, and knowledge distillation, highlighting relevant technologies. In Section III, we present the overall architecture of a 3D object detection framework with reflectance prediction-based knowledge distillation and elaborate on its key components. Section IV demonstrates the effectiveness of our method on the KITTI dataset and Waymo Open Dataset and includes ablation studies to evaluate the effect of each component. Finally, Section V concludes this paper.
@@ II Related Work
In 3D object detection for raw point clouds, existing methods can be sorted into two classes according to the data format of network input: point-based [14, 15, 16, 17, 18, 19, 20, 21] and voxel-based [22, 23, 24, 25, 26, 27, 28, 29, 30, 31] methods. Point-based detectors extract hierarchical point-wise features by various sampling and aggregation methods to predict 3D objects. In contrast, voxel-based detectors utilize 3D sparse convolution to extract voxel features, and these features are projected onto the Bird’s Eye View (BEV) for 3D object detection. In 3D object detection for compressed point clouds, those generated by range image-based [32, 33, 34, 35] and voxel-based [8, 9] compression frameworks are input into compressed point-cloud detectors for recognition and localization. These methods generally rely on raw point clouds for pre-training 3D detectors and assess detection performance on lossy compressed point clouds across various code rates using the pre-trained models, as shown in Fig. 1(a) and 1(b). However, they lack effective robustness measures against information loss and reconstruction errors caused by lossy point cloud compression, leading to a significant decline in detection accuracy.
For robust 3D object detection in compressed point clouds, this paper addresses three key aspects. First, voxelized representation of compressed point clouds helps mitigate the negative effects of lossy compression. Thus, we develop robust compressed point-cloud detectors on voxel-based detection frameworks. Second, pre-training 3D detectors solely on raw point clouds is insufficient for adapting to the characteristics of compressed point clouds. To improve detection performance, our 3D detectors are trained on both raw and compressed point clouds. Finally, with a two-source joint training scheme, we use voxel-based two-stage detectors like PV-RCNN [22] and Voxel-RCNN [23] as backbone networks and transfer detection knowledge of the first-stage proposals from raw to compressed point-cloud detectors, thereby enhancing detection robustness for compressed point clouds.
@@ II-B Point Cloud Compression
LiDAR point cloud compression methods can be categorized into three classes based on data representation: point-based [36, 37, 38], image-based [32, 33, 34, 39, 40, 35, 41, 42, 43], and voxel-based [5, 6, 7, 8, 9, 10, 11, 12, 13, 44] methods. Point-based methods utilize flexible point sampling and feature extraction to generate point-wise features, achieving point cloud compression and reconstruction. However, these methods are confronted with uneven 3D object loss and noise-point interference due to the lack of spatial structure and context information during lossy compression. Image-based methods project point clouds onto 2D images for quantization and encoding, but they introduce more severe three-dimensional spatial distortion and information loss compared to voxel-based methods. By analyzing octree or multi-scale sparse tensors, voxel-based methods enhance compression performance via context feature learning, multi-scale joint optimization, and redundant encoding reduction. Among these compression methods, only G-PCC [5] and Unicorn [7] incorporate reflectance encoding modules, as reflectance encoding presents challenges in structural design, end-to-end optimization, and data-flow overhead. To solve this problem, we design an RP module at the receiver to supplement reflectance information for detection training on non-reflectance compressed point clouds generated by voxel-based methods. Our approach effectively alleviates the design complexity of LiDAR point cloud compression and provides an end-to-end optimization strategy for reflectance prediction at the receiver.
@@ II-C Knowledge Distillation
Knowledge distillation methods for 3D object detection can be divided into three categories according to their research objectives: model compression distillation [45, 46, 47, 48, 49, 50], detection improvement distillation [51, 52], and weak-data robustness distillation [53, 54, 55]. In model compression distillation, SparseKD [45] introduces response-based constraints using Logit and Label KD to guide lightweight student models, achieving accurate and efficient 3D object detection. In detection improvement distillation, SPNet [51] assigns category values to the interior points of ground truth (GT) boxes, while X-Ray [52] addresses object point loss caused by sparsity and occlusion utilizing tracking information from consecutive-frame point clouds. Both methods augment data representations with strong semantic priors, providing valuable supervisory knowledge for distillation. In weak-data robustness distillation, student models employ low-quality point clouds from rainy conditions [55] or low-resolution sensor data [53, 54], learning complete and accurate knowledge from raw point clouds to improve detection robustness in weak data scenarios. Inspired by these weak-data robustness distillation methods, our RPKD framework treats non-reflectance compressed point clouds involved in vehicle networking data transmission as weak data for student models and raw point clouds as guiding data for teacher models, aiming to develop robust compressed point-cloud detectors.
@@ III-A Overview of our Framework
During point cloud compression, reflectance encoding faces significant challenges in both end-to-end optimization and data-flow overhead. According to the separable compression mechanism of G-PCC [5], existing methods typically disregard reflectance compression and focus solely on optimizing geometric coordinate encoding to achieve efficient transmission and accurate detection. As a result, decoded compressed point clouds often lack reflectance information and suffer from point number reduction and reconstruction errors, which limit 3D object detection performance. To address this issue, we propose a 3D object detection framework with reflectance prediction-based knowledge distillation (RPKD), as shown in Fig. 3, to obtain more robust and precise 3D detectors against non-reflectance compressed point clouds.
First, our reflectance cross-match (RCM) module correlates lossy compressed points with neighboring raw point-cloud voxels. The nearest voxel’s average reflectance is assigned as the reflectance label for the current compressed point. Similarly, the reflectance inter-match (RIM) module assigns a voxel’s average reflectance as the reflectance label for corresponding raw points. Next, the proposed reflectance prediction (RP) module generates point-wise geometric features using 3D sparse convolution and voxel set extraction to predict reflectance. The predicted point-wise reflectance is then incorporated into corresponding non-reflectance voxels for subsequent detection tasks. Finally, we treat non-reflectance compressed point clouds as weak data and introduce two response-based knowledge distillation constraints: reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD). Specifically, RKD transfers reflectance prediction knowledge from raw point clouds via the RP module, while DKD conveys first-stage proposal knowledge from the two-stage 3D detector for raw point clouds. In what follows, we detail the proposed framework.
@@ III-B Reflectance Label Generation
Due to point number reduction and reconstruction errors from lossy compression, compressed points do not correspond one-to-one with raw points. For compressed point reflectance prediction, we introduce an RCM module based on spatial relationships to assign reflectance labels from raw point-cloud voxels to corresponding compressed points. These labels, used as prior knowledge for reflectance prediction, are generated frame by frame during detection training and do not need to be compressed or transmitted. Additionally, to transfer effective knowledge during reflectance distillation, we also design a similar RIM module to acquire reflectance labels for raw points. The principles of these modules are outlined below.
@@
Reflectance Cross-match Module. For geometric coordinate encoding, voxel-based compression methods [6, 9] are often employed to obtain non-reflectance compressed point clouds . Before training, with each point’s reflectance assigned to 0, the compressed point clouds fed into our detection networks are . Raw point clouds are represented as . These point clouds are voxelized to the same scale, generating compressed point-cloud voxels and raw point-cloud voxels . When using , and as inputs, the principle of our RCM module can be described as follows.
@@
As depicted in Fig. 4, there are two spatial matching strategies between compressed points and raw point-cloud voxels: center matching and surrounding matching. Specifically, the voxel corresponding to compressed point coordinates is denoted as , and the surrounding non-empty voxels within a 1-voxel matching range are . The initial reflectance labels for compressed points are defined as:
@@
where represents the average reflectance of in center matching, and is the nearest voxel’s average reflectance in surrounding matching. However, when multiple compressed points exist within the same compressed point-cloud voxel , their reflectance labels may come from different raw point-cloud voxels, resulting in significant numerical discrepancies that negatively impact compressed point reflectance prediction. To address this issue, we further calculate the average reflectance labels for all points within each compressed point-cloud voxel . The final reflectance labels for compressed points are thus expressed as:
@@
where denotes the reflectance label for each point within the current compressed point-cloud voxel , and represents the averaging operation applied to these reflectance labels.
@@
Reflectance Inner-match Module. In the RPKD framework, it is crucial to establish suitable reflectance prediction knowledge from raw point clouds to guide the prediction process for compressed point-cloud reflectance. Specifically, we treat the reflectance predictions for compressed and raw point clouds as a student-teacher pair, where both sets of labels need to maintain spatial correspondence and information similarity. In the RCM module, the reflectance labels for compressed points are derived from the corresponding raw point-cloud voxels. Therefore, our RIM module avoids directly using raw point reflectance and instead assigns the average reflectance of raw point-cloud voxels to corresponding points as their reflectance labels. Since raw points exactly correspond to their voxels, our RIM module employs center matching. With raw point coordinates and voxels as inputs, the raw point reflectance labels are computed by the following formula:
where is the voxel corresponding to raw point coordinates , and denotes the average reflectance of this voxel.

### v2

@@ Abstract
Regarding intelligent transportation systems, low-bitrate transmission via lossy point cloud compression is vital for facilitating real-time collaborative perception among connected agents, such as vehicles and infrastructures, under restricted bandwidth. In existing compression transmission systems, the sender lossily compresses point coordinates and reflectance to generate a transmission code stream, which faces transmission burdens from reflectance encoding and limited detection robustness due to information loss. To address these issues, this paper proposes a 3D object detection framework with reflectance prediction-based knowledge distillation (RPKD). We compress point coordinates while discarding reflectance during low-bitrate transmission, and feed the decoded non-reflectance compressed point clouds into a student detector. The discarded reflectance is then reconstructed by a geometry-based reflectance prediction (RP) module within the student detector for precise detection. A teacher detector with the same structure as the student detector is designed for performing reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD) from raw to compressed point clouds. Our cross-source distillation training strategy (CDTS) equips the student detector with robustness to low-quality compressed data while preserving the accuracy benefits of raw data through transferred distillation knowledge. Experimental results on the KITTI and DAIR-V2X-V datasets demonstrate that our method can boost detection accuracy for compressed point clouds across multiple code rates. We will release the code publicly at https://github.com/HaoJing-SX/RPKD.
I Introduction
@@
LiDAR point clouds, composed of sparse points with both geometric (coordinates) and attribute (reflectance) information about object surfaces, are vital for predicting the categories, positions, and bounding boxes of 3D objects. In vehicle-to-everything (V2X) cooperative perception, receivers need to rapidly acquire LiDAR point clouds captured by neighboring senders to enable an expanded perception range and accurate object detection [1, 2, 3, 4]. Raw point clouds are extremely voluminous and require a high transmission bandwidth, making it difficult to realize real-time cooperative perception. To this end, point cloud compression reduces the amount of transmitted data, improving data acquisition efficiency at the receiver. However, compressed point clouds generated by existing methods often lose some geometric coordinates and reflectance information, leading to suboptimal performance in 3D object detection. Therefore, enhancing the 3D detector’s robustness for compressed point clouds is of great research significance.
@@
Existing 3D object detection methods in compressed point clouds can be classified into two categories based on their reflectance processing approaches: 3D object detection with [5, 6, 7] or without reflectance encoding [8, 9, 10, 11, 12, 13]. For 3D object detection with reflectance encoding, both G-PCC [5] and Unicorn [6, 7] compressed voxelized point-cloud coordinates together with reflectance, generating geometry and attribute bitstreams for reconstruction and detection tasks, as shown in Fig. 1(a). However, the attribute bitstream in these methods introduces a transmission burden that multiplies with the number of connected vehicles and infrastructures. To solve this issue, in 3D object detection without reflectance encoding, existing compression methods based on voxelized point clouds and deep learning [8, 9, 10, 11, 12, 13] typically focused on compressing geometric coordinates while discarding reflectance to elevate transmission efficiency, as presented in Fig. 1(b). However, lossy compression by these methods causes substantial information loss in low-bitrate transmission, severely impacting 3D object detection accuracy. To address these challenges, this paper adopts the voxel-based lossy geometric compression method PCC-S [9] to maximize bandwidth reduction and systematically investigates training strategies for non-reflectance compressed point clouds to develop a robust detector.
The existing raw-data single-source training strategy (STS-R) [8, 14] pretrained detection models on raw data and evaluated them on low-quality compressed data, resulting in limited robustness. To handle the low quality of compressed point clouds, we adopt a compressed-data single-source training strategy (STS-C), training and evaluating at each compression rate to improve detection performance. In 3D object detection knowledge distillation, FD [15] transferred classification, regression, and feature knowledge from high-resolution to low-resolution data, thereby enhancing detection accuracy on low-resolution inputs. Building on this idea, we propose a cross-source distillation training strategy (CDTS), in which raw and compressed point-cloud models form a teacher-student model pair, and response-based detection knowledge distillation (DKD) from SparseKD [16] is applied to transfer accurate detection knowledge, as illustrated in Fig. 1(c).
Using PV-RCNN as the backbone, Fig. 2 reports performance for different training strategies and detection methods on non-reflectance compressed data. SparseKD-PV (CDTS) outperforms PV-RCNN (STS-C) by an average of 1.11 mAP across various data types, yet still lags behind the PV-RCNN baseline on raw data. In general, LiDAR reflectance offers critical cues for 3D object recognition, thereby enhancing detection reliability. For SparseKD-PV (CDTS), however, discarding reflectance to improve transmission efficiency remains a key factor contributing to the suboptimal performance. To overcome this limitation, we introduce a dedicated reflectance prediction (RP) module at the receiver and employ knowledge distillation techniques to refine reflectance estimation, ultimately boosting detection performance on compressed point clouds.
@@
Incorporating the previous analysis, we propose a 3D object detection framework with reflectance prediction-based knowledge distillation (RPKD) to improve detection accuracy for non-reflectance compressed point clouds, as depicted in Fig. 1(d). Since compressed and raw point clouds lack direct point-to-point correspondence, we design a reflectance cross-match (RCM) module to assign reflectance labels from raw point-cloud voxels to compressed points. To construct suitable knowledge for reflectance distillation, we introduce a comparable reflectance inter-match (RIM) module to generate reflectance labels for the raw point-cloud detector. Besides, the proposed reflectance prediction (RP) module comprehensively extracts geometric features of voxelized point clouds, achieving reflectance reconstruction at the receiver. Our cross-source distillation training strategy (CDTS) utilizes response-based reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD) to transfer guided knowledge from raw to compressed point-cloud detectors, enabling more accurate reflectance reconstruction and improved 3D object detection. As shown in Fig. 2, RPKD-PV (CDTS) significantly raises the mAP values across various data types, demonstrating the effectiveness of our method for non-reflectance compressed point clouds.
@@
The proposed CDTS transfers supervisory knowledge from raw to compressed point-cloud detectors using two response-based distillation constraints, RKD and DKD, further enhancing detection robustness for low-quality compressed data.
The remainder of the paper is organized as follows. Section II reviews existing studies in terms of 3D object detection, point cloud compression, and knowledge distillation, highlighting relevant technologies. In Section III, we present the overall architecture of a 3D object detection framework with reflectance prediction-based knowledge distillation and elaborate on its key components. Section IV demonstrates the effectiveness of our method on the KITTI and DAIR-V2X-V datasets, and includes ablation studies to assess the impact of each component. Finally, Section V concludes this paper.
@@ II Related Work
In 3D object detection for raw point clouds, existing methods can be sorted into two classes according to the data format of network input: point-based [17, 18, 19, 20, 21, 22] and voxel-based [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] methods. Point-based detectors extract hierarchical point-wise features by various sampling and aggregation methods to predict 3D objects. In contrast, voxel-based detectors utilize 3D sparse convolution to extract voxel features, and these features are projected onto the bird’s eye view (BEV) for 3D object detection. In 3D object detection for compressed point clouds, those generated by range image-based [35, 14, 36, 37] and voxel-based [5, 6, 7, 8, 9, 10, 11, 12, 13] compression frameworks were input into detectors for recognition and localization. These methods typically rely on raw point clouds for pretraining 3D detectors and evaluate detection performance on lossy compressed point clouds at various code rates using the pretrained models. However, the raw-data single-source training strategy lacks adaptability to low-quality compressed data, resulting in a significant decline in detection accuracy. To address this, we propose a novel CDTS that improves detector robustness to low-quality compressed data while inheriting the accuracy of raw data.
@@ II-B Point Cloud Compression
LiDAR point cloud compression methods can be categorized into three classes based on data representation: point-based [38, 39, 40], image-based [35, 14, 36, 41, 42, 37, 43, 44, 45], and voxel-based [5, 6, 7, 8, 9, 10, 11, 12, 13, 46] methods. Point-based methods utilize flexible point sampling and feature extraction to generate point-wise features, achieving point cloud compression and reconstruction. However, these methods are confronted with uneven 3D object loss and noise-point interference due to the lack of spatial structure and context information during lossy compression. Image-based methods project point clouds onto 2D images for quantization and encoding, but they introduce more severe three-dimensional spatial distortion and reconstruction errors compared to voxel-based methods. By analyzing octree or multi-scale sparse tensors, voxel-based methods enhance compression performance via context feature learning, multi-scale joint optimization, and redundant encoding reduction. Among these compression methods, only G-PCC [5] and Unicorn [7] incorporated reflectance encoding modules, as reflectance encoding presents challenges in structural design, end-to-end optimization, and data-flow overhead. To solve this problem, we design an RP module at the receiver to supplement reflectance information for detection training on non-reflectance compressed point clouds generated by voxel-based methods. Our approach effectively alleviates the design complexity of LiDAR point cloud compression and provides an end-to-end optimization strategy for reflectance prediction at the receiver.
@@ II-C Knowledge Distillation
Knowledge distillation methods for 3D object detection can be divided into three categories according to their research objectives: model compression distillation [16, 47, 48, 49, 50, 51], detection improvement distillation [52, 53], and weak-data robustness distillation [54, 15, 55]. In model compression distillation, SparseKD [16] introduced response-based constraints through Logit and Label KD to guide lightweight student models, achieving accurate and efficient 3D object detection. In detection improvement distillation, SPNet [52] assigned category values to the interior points of ground truth (GT) boxes, while X-Ray [53] addressed object point loss caused by sparsity and occlusion utilizing tracking information from consecutive-frame point clouds. Both methods augment data representations with strong semantic priors, providing valuable supervisory knowledge for distillation. In weak-data robustness distillation, student models employed low-quality point clouds from rainy conditions [55] or low-resolution sensor data [54, 15], learning complete and accurate knowledge from raw point clouds to improve detection robustness in weak-data scenarios. Inspired by weak-data robustness distillation methods, our RPKD framework treats non-reflectance compressed point clouds in vehicle networking data transmission as weak data for student models, with raw point clouds guiding teacher models, aiming to develop robust compressed point-cloud detectors.
II-D V2X Collaborative Perception
To overcome single-vehicle perception limitations, V2X collaborative perception [3, 4, 56, 57, 58] develops early, intermediate, and late fusion strategies based on shared information among connected agents. These strategies correspond to the transmission of raw data, intermediate layer features, and prediction results, respectively. In early fusion with LiDAR [1, 2], raw point clouds offer precise scene information and are easily registered, but they require significant transmission bandwidth. Cooper [1] mitigated the bandwidth demand for mutual transmission between two vehicles by extracting key point clouds using RoIs. However, this approach still faces challenges in scaling to large V2X collaborative scenarios. To address these issues, this paper pioneeringly introduces voxel-based lossy geometric compression into collaborative perception, effectively balancing low bandwidth and high accuracy. Regarding datasets, DAIR-V2X [58] is more mature and was released earlier than V2X-Radar [56]. For detection tasks, we focus on single-view 3D detection in DAIR-V2X-V and evaluate performance on compressed data across various object categories, thus laying the foundation for future collaborative detection.
@@ III-A Overview of our Framework
In collaborative perception with LiDAR point cloud compression, reflectance encoding imposes a linearly increasing transmission burden as the number of connected agents grows. According to the separable compression mechanism of G-PCC [5], existing methods typically disregard reflectance compression and focus solely on optimizing geometric coordinate encoding to achieve efficient transmission and accurate detection. As a result, decoded compressed point clouds often lack reflectance information and suffer from point number reduction and reconstruction errors, which limit 3D object detection performance. To address this issue, we propose a 3D object detection framework with reflectance prediction-based knowledge distillation (RPKD), as shown in Fig. 3, to obtain more robust and precise 3D detectors against non-reflectance compressed point clouds.
First, our reflectance cross-match (RCM) module correlates lossy compressed points with neighboring raw point-cloud voxels. Average reflectance of the nearest voxel is assigned as the reflectance label for the current compressed point. Similarly, the reflectance inter-match (RIM) module assigns the average reflectance of a voxel as the label for corresponding raw points. Next, the proposed reflectance prediction (RP) module generates point-wise geometric features using 3D sparse convolution and voxel set extraction to predict reflectance. The predicted reflectance is then integrated into the corresponding non-reflectance voxels for subsequent detection tasks. Finally, our cross-source distillation training strategy (CDTS) transfers guiding knowledge from raw to compressed point-cloud detectors through two response-based distillation constraints. Specifically, reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD) concentrate on the RP module’s reflectance knowledge and the first-stage proposal knowledge, respectively. In what follows, we detail the proposed framework.
@@ III-B Reflectance Label Generation
Due to point number reduction and reconstruction errors from voxel-based lossy compression, compressed points do not correspond one-to-one with raw points. For compressed point reflectance prediction, we introduce an RCM module based on spatial relationships to assign reflectance labels from raw point-cloud voxels to corresponding compressed points. These labels, used as prior knowledge for reflectance prediction, are generated frame by frame during detection training and do not need to be compressed or transmitted. Additionally, to transfer effective knowledge during reflectance distillation, we also design a similar RIM module to acquire reflectance labels for raw points. The principles of these modules are outlined below.
@@
Reflectance Cross-match Module. For geometric coordinate encoding, voxel-based compression methods [6, 9] are often employed to obtain non-reflectance compressed point clouds. With each compressed point’s reflectance assigned to zero, raw and compressed point clouds fed into our detection networks are represented as and . These point clouds are voxelized to the same scale, generating raw point-cloud voxels and compressed point-cloud voxels . When using the compressed point coordinates and the voxels of raw and compressed point clouds as inputs, the principle of our RCM module can be described as follows.
@@
As depicted in Fig. 4, there are two spatial matching strategies between compressed points and raw point-cloud voxels: center matching and surrounding matching. Specifically, the center voxel corresponding to compressed point coordinates is denoted as , and the surrounding non-empty voxels within a 1-voxel matching range are . The initial reflectance labels for compressed points are defined as:
@@
where represents the average reflectance of a center voxel, and is the nearest voxel’s average reflectance in surrounding matching. However, when multiple compressed points exist within a compressed point-cloud voxel, their reflectance labels may come from different raw point-cloud voxels, resulting in significant numerical discrepancies that negatively impact compressed point reflectance prediction. To address this issue, we further calculate the average reflectance labels for all points within each compressed point-cloud voxel. The final reflectance labels for compressed points are thus expressed as:
@@
where denotes the reflectance label for each point within the current compressed point-cloud voxel , and represents the averaging operation applied to these reflectance labels, with and denoted as the corresponding numbers of compressed points and voxels, respectively.
@@
Reflectance Inner-match Module. In the RPKD framework, it is crucial to establish suitable reflectance knowledge from raw point clouds to guide the prediction process for compressed point-cloud reflectance. Specifically, we treat the reflectance predictions for raw and compressed point clouds as a teacher-student pair, where both sets of labels need to maintain spatial correspondence and information similarity. In the RCM module, the reflectance labels for compressed points are derived from the corresponding raw point-cloud voxels. Therefore, our RIM module avoids directly using raw point reflectance and instead assigns the average reflectance of raw point-cloud voxels to corresponding points. Since raw points correspond exactly to their voxels, the RIM module uses center matching to obtain the raw point reflectance labels .
These pluggable matching modules are designed for non-reflectance compressed point clouds, which are obtained through voxel-based geometric compression methods such as PCC-S. When compressed point clouds inherently include reflectance, the reflectance labels are directly available.